OpenStudy (anonymous):

determine the general solution: sin 2x=-1/2

myininaya (myininaya):

remember sin(7pi/6)=-1/2 and sin(11pi/6)=-1/2 so sin(7pi/6+2npi)=-1/2 and sin(11pi/6+2npi)=-1/2 for n=0,1,2,3,... so 2x=7pi/6+2npi and 2x=11pi/6+2npi solve both of these for x and you are done

OpenStudy (anonymous):

let 2x = alpha or beta or theta sin(a)=-1/2 a=30 degrees right ( taking Sin^-1 * -1/2) we know using unit circle that sin is negative in 3rd and 4th quad one angle will be 30 and other will be 330. as a =2x 2x=30 2x=330 x=15 and x=165

OpenStudy (anonymous):

there you have solutions

myininaya (myininaya):

sin30 is not -1/2

OpenStudy (anonymous):

agreed

myininaya (myininaya):

i think he knew that because he did say look in 3rd and 4th quad he just made mistake

OpenStudy (anonymous):

yes,seems to be

myininaya (myininaya):

uzma am i your fan i should be your fan let me look

OpenStudy (anonymous):

:P

myininaya (myininaya):

yep i am

OpenStudy (anonymous):

same here;)

OpenStudy (anonymous):

i dont know how did i miss it earlier

OpenStudy (anonymous):

hopefully u were not making a fun of me :P

myininaya (myininaya):

lol you will never know i go to sleep now goodnight

OpenStudy (anonymous):

sorry that was a typing error. buh i did mention it

OpenStudy (anonymous):

-> sin 2x=-1/2 -> sin 2x=sin(-pi/6) -> 2x=n(pi)+(pi/6)(-1)^n where n is integer (this is a formula,if sin x=sin a ,and ( -pi/2) <= (a ) <=(pi/2) then x= n(pi) + ( -1) to the power n multiplied by (a) ) -> x=(1/2) *( n(pi)+(pi/6)(-1)^n ) where n is integer .....for values of x just put values of n= ... -3,-2,-1,0,1,2,3....